# B3C: A Minimalist Approach to Offline Multi-Agent Reinforcement Learning

### v2

@@ Abstract
Overestimation arising from selecting unseen actions during policy evaluation is a major challenge in offline reinforcement learning (RL). A minimalist approach in the single-agent setting—adding behavior cloning (BC) regularization to existing online RL algorithms—has been shown to be effective; however, this approach is understudied in multi-agent settings. In particular, overestimation becomes worse in multi-agent settings due to the presence of multiple actions, resulting in the BC regularization-based approach easily suffering from either over-regularization or critic divergence. To address this, we propose a simple yet effective method, Behavior Cloning regularization with Critic Clipping (B3C), which clips the target critic value in policy evaluation based on the maximum return in the dataset and pushes the limit of the weight on the RL objective over BC regularization, thereby improving performance. Additionally, we leverage existing value factorization techniques, particularly non-linear factorization, which is understudied in offline settings. Integrated with non-linear value factorization, B3C outperforms state-of-the-art algorithms on various offline multi-agent benchmarks.
Machine Learning, ICML
1 Introduction
@@
Reinforcement learning (RL) trains agents to maximize cumulative rewards through interaction with an environment in an online manner; however, such interaction can be time-consuming due to the sample inefficiency of RL and costly in safety-critical environments due to inherent risks. Furthermore, only existing datasets are available for use in practice. These challenges have motivated a shift toward offline RL, which trains agents using pre-collected datasets without requiring interaction.
@@
A major challenge of offline RL is that bootstrapping from unseen actions in policy evaluation induces extrapolation error, leading to accumulated errors (Fujimoto & Gu, 2021), due to the inability to generate data. This issue consequently causes overestimation of value functions and adversely affects policy improvement. To address this, a variety of techniques, such as conservative value estimation (Kumar et al., 2020) and modeling behavior policy (Fujimoto et al., 2019), have been introduced. However, these approaches increase algorithmic complexity, hinder reproducibility, and amplify hyperparameter sensitivity. This motivates minimalist approaches, which introduce minimal changes to existing RL algorithms (Fujimoto & Gu, 2021; Tarasov et al., 2024). One representative example is TD3+BC, which integrates behavior cloning (BC) regularization into an existing online RL algorithm called TD3 (Fujimoto & Gu, 2021). This simple approach does not require any additional components and introduces only one more hyperparameter for regularization, but it surprisingly achieves SOTA performance. This line of research—minimalist approaches—has been extensively investigated in single-agent settings (Fujimoto & Gu, 2021; Tarasov et al., 2024), but it has barely been explored in multi-agent settings, which entail more algorithmic complexity due to the presence of multiple agents.
@@
To address the aforementioned limitations, we propose a simple yet effective technique for offline multi-agent RL: Behavior Cloning regularization with Critic Clipping (B3C). The proposed method alleviates overestimation by clipping the critic value during policy evaluation, using the maximum return from the given dataset, in addition to BC regularization. Critic clipping enables the use of a higher weight for the RL objective relative to BC regularization, resulting in improved performance. Furthermore, we integrate B3C with an existing online multi-agent RL algorithm called FACMAC (Peng et al., 2021), which utilizes factored critics and value factorization techniques, resulting in FACMAC+B3C. Notably, we provide a design choice for value factorization in the offline setting, empirically demonstrating that non-monotonic factorization performs better than monotonic and linear factorization, which contrasts with findings in the online setting (Peng et al., 2021).
@@
Despite its simplicity, the effectiveness of FACMAC+B3C is empirically demonstrated across a variety of offline multi-agent RL benchmarks, including multi-agent Mujoco and particle environments, with various types of dataset. We provide a thorough analysis demonstrating how the proposed method outperforms the baselines in terms of performance.
2 Background and Related Works
2.1 Multi-Agent RL
@@
We consider fully cooperative multi-agent tasks, which can be modeled as an N-agent decentralized partially observable MDP (Dec-POMDP) (Oliehoek, 2012). At each timestep $t$, each agent selects its own action, $a_{t}^{i}$, based on its local information such as the partial observation $o_{t}^{i}$. This forms a joint action $\bm{a}_{t}$ yielding the next state $s_{t+1}$, the joint observations $\bm{o}_{t+1}=(o^{1}_{t+1},\cdots,o^{N}_{t+1})$, and a shared reward $r_{t}$. The goal is to find the optimal joint policy that maximizes the team return, $\mathbb{E}[\sum_{l=0}^{\infty}\gamma^{l}r_{l}]$. For this, recent multi-agent RL algorithms adopt a framework of centralized training with decentralized execution (CTDE) (Wang et al., 2020; Zhang et al., 2021; Jeon et al., 2022; Kim et al., 2023), where individual policies execute actions based on local information, e.g., Agent $i$’s action-observation history $\tau^{i}$, but are trained with global information such as the state.
@@
FACMAC (Peng et al., 2021) is a representative multi-agent RL algorithm that trains a centralized but factored critic, which is decomposed into a nonlinear combination of individual critics, and uses it to train decentralized deterministic policies.The key idea is factorizing the centralized critic into individual critics via a non-linear function, where the centralized critic is decomposed as $Q_{jt}(s,\bm{\tau},\textbf{a})=f_{mixer}(s,Q^{1}(\tau^{1},a^{1}),\cdots,Q^{N}(\tau^{N},a^{N}))$, where $f_{mixer}$ is a mixer network that encodes state information and learns the weights of individual critics, and $Q^{i}(\tau^{i},a^{i})$ is the individual critic of Agent $i$. FACMAC considers two value factorization methods: (a) monotonic factorization (mono), which constrains the parameters of $f_{mixer}$ to enforce $\partial Q_{jt}/\partial Q^{i}\geq 0$, as in QMIX (Rashid et al., 2020); and (b) non-monotonic factorization (non-mono), which removes the constraint from (a) to increase representational capacity. In addition to these two methods (c) linear factorization (Sunehag et al., 2017) (vdn) was also considered. It was shown that non-mono and mono approaches are complementary: non-mono performs better on tasks requiring greater representational capacity, while mono outperforms on others. However, vdn underperforms both and performs drastically worse, particularly in multi-agent Mujoco environment.
@@
$\mathcal{L}_{\text{FACMAC}}$ $(\bm{\theta_{Q}})=\mathbb{E}_{\mathcal{D}}\left[\Big{(}y^{jt}-Q_{jt}(s,\bm{\tau},\bm{a};\bm{\theta_{Q}})\Big{)}^{2}\right],$
$\mbox{where}~{}y^{jt}=r+\gamma Q_{jt}(s^{\prime},\bm{\tau^{\prime}},\bm{\pi}(\bm{\tau^{\prime}});\bm{\theta_{Q}}^{-}),$ (1)
$\mathcal{D}$ is the replay buffer, $\bm{\pi}$ is the deterministic joint policy, and $\bm{\theta_{Q}^{\prime}}$ are the parameters of the target critic networks. Based on this centralized critic, the decentralized policies, parameterized by $\bm{\theta}_{\bm{\pi}}$, are trained using the deterministic policy gradient (Silver et al., 2014) where the loss function is written as $\mathcal{L}_{\text{FACMAC}}(\bm{\theta}_{\bm{\pi}})=$
@@
where $\theta_{\pi}^{i}$ is Agent $i$’s actor parameter.
A main benefit of FACMAC over prior works such as MADDPG (Lowe et al., 2017) is leveraging non-linear factorization. However, in offline settings, value factorization techniques remain underexplored—recent research still relies on linear factorization (Yang et al., 2021; Wang et al., 2024) or omits it by using a non-factored critic. In this paper, to investigate value factorization for offline settings, we adopt FACMAC to build upon the recent achievements.
2.2 Offline RL
Offline RL aims to learn a policy that maximizes the expected return from a fixed dataset $\mathcal{D}$, consisting of trajectories generated by arbitrary behavior policies, without additional environment interactions (Kumar et al., 2020; Kostrikov et al., 2022; Kim et al., 2024a). This inability to generate additional data introduces a major challenge in offline RL—extrapolation error in policy evaluation, i.e., the target value in the Bellman equation becomes inaccurate if actions from the learning policy are not included in the dataset. This extrapolation error often leads to overestimation and degrades performance. To address this challenge, several approaches have been proposed, such as behavior-constrained policy optimization, which regularizes the policy to stay close to the behavior policy (Wu et al., 2019; Fujimoto & Gu, 2021), and conservative value estimation (Kumar et al., 2020; Kostrikov et al., 2021), which lower-bounds the value function by penalizing the values of unseen actions. These approaches have demonstrated effectiveness; however, they often require significant modifications or additional components to existing RL algorithms, such as generative models. These additions introduce more hyperparameters, which can affect performance and impact stability and reproducibility. This necessitates an offline RL algorithm with minimal changes from the current online RL algorithms. As an example, (Fujimoto & Gu, 2021) proposed TD3 with Behavior Cloning (TD3+BC), which adds a BC loss function as a regularizer on top of the TD3 loss (Fujimoto et al., 2018) to update the policy. This encourages the action generated by the learning policy to be the same as the action in the dataset. The policy of TD3+BC is updated to minimize the following loss function.
@@ 2.2 Offline RL
$\mbox{where}~{}~{}~{}\bm{w}=\frac{\alpha}{\frac{1}{N}\sum|Q(s,\pi(s))|}$ (3)
@@
where $\alpha$ is a hyperparameter that controls the balance between RL and BC. TD3+BC has been widely used due to its simplicity and performance comparable to SOTA algorithms (Kim et al., 2024b).
2.3 Offline Multi-Agent RL
@@
A natural approach to offline multi-agent RL is to extend the previously mentioned single-agent offline RL to the multi-agent setting. However, naively extending these methods has been shown to be insufficient (Pan et al., 2022; Shao et al., 2024). For example, naive conservative value estimation for the joint policy can lead to excessively large penalties, significantly degrading performance as the number of agents increases (Shao et al., 2024). Furthermore, the problem of extrapolation error worsens as the size of the action space grows with the number of agents (Yang et al., 2021), making it essential to balance conservatism and regularization with extrapolation error. To mitigate this problem, CFCQL (Shao et al., 2024) introduces counterfactual regularization to each agent individually to alleviate excessive conservatism. OMAR (Pan et al., 2022) introduces zeroth-order optimization to address policies getting stuck in poor local optima under conservative value function training. Note that both OMAR and CFCQL are based on conservative value estimation, which penalizes value functions for unseen action spaces, thereby forcing them to underestimate. OMIGA (Wang et al., 2024) introduces a framework that converts global regularization into implicit local-level regularization using linear value decomposition, while employing in-sampling techniques to avoid querying unseen actions. MADIFF (Zhu et al., 2023) presents a diffusion-based generative approach for multi-agent systems, enabling decentralized policy development through effective teammate modeling.
@@
Although the aforementioned offline multi-agent RL algorithms perform reasonably, they still require significant modifications from existing multi-agent RL algorithms, as described in single-agent settings in Sec. 2.2, with even more adjustments needed due to the presence of multiple agents. Additionally, they have not leveraged the currently existing value factorization techniques, which are central to the recent success of multi-agent RL. Therefore, inspired by (Fujimoto & Gu, 2021), we aim to develop a minimalist approach that minimally modifies an existing multi-agent RL algorithm, focusing on simplicity and practicality while incorporating existing multi-agent RL techniques.
3 Methodology
@@
To achieve the aforementioned goal, we focus on a BC regularization-based approach. In multi-agent settings, the presence of multiple agents amplifies the issue of overestimation, as mentioned in Sec.2.3. Consequently, the BC regularization-based approach is susceptible to either over-regularization or unstable critic learning, which often stems from insufficient regularization. We propose a simple yet effective regularization method called Behavior Cloning with Critic Clipping (B3C), which prevents over-regularization while ensuring stability, allowing a focus on the RL objective and thereby significantly improving performance. Additionally, we explore the application of value factorization techniques with B3C, which remains understudied.
@@
Figure 1: Test return with respect to dataset quality (average return of the dataset) in the Cooperative Navigation (CN) and World environments, reported in Shao et al. (2024). Performance close to $y=x$ indicates a strong dependency on data quality.
3.1 Revisiting Behavior Cloning: Over-regularization
We begin by revisiting BC regularization in offline multi-agent RL. Building on the success of BC regularization in the single-agent setting, recent studies on offline multi-agent RL (Pan et al., 2022; Shao et al., 2024) have introduced an extension of TD3+BC, called MA-TD3+BC, as a baseline by adapting Eq. 2.2 to the multi-agent setting. Although MA-TD3+BC achieves reasonable performance, it is less effective than the proposed methods, particularly when applied to low-quality datasets. Fig. 1 illustrates the test returns of OMAR (Pan et al., 2022), CFCQL (Shao et al., 2024), and MA-TD3+BC with respect to the average return of the dataset, reported in (Shao et al., 2024). MA-TD3+BC achieves performance close to dataset-level quality, but with a small margin. In contrast, OMAR and CFCQL show significant improvements over dataset quality, particularly for low-quality datasets. This observation, that MA-TD3+BC heavily depends on data quality, indicates that it is over-regularized by behavior cloning.

### v3

@@ Abstract.
Overestimation arising from selecting unseen actions during policy evaluation is a major challenge in offline reinforcement learning (RL). A minimalist approach in the single-agent setting—adding behavior cloning (BC) regularization to existing online RL algorithms—has been shown to be effective in terms of achieving competitive performance with minimal modification to online algorithms; however, this approach is understudied in multi-agent settings. In particular, overestimation becomes worse in multi-agent settings due to the presence of multiple actions, resulting in the BC regularization-based approach easily suffering from either over-regularization or critic divergence. To address this, we propose a simple yet effective method, Behavior Cloning regularization with Critic Clipping (B3C), which clips the target critic value in policy evaluation based on the maximum return in the dataset and pushes the limit of the weight on the RL objective over BC regularization, thereby demonstrating superior performance across benchmarks. Additionally, we leverage existing value factorization techniques, particularly non-linear factorization, which is understudied in offline settings. Integrated with non-linear value factorization, B3C outperforms state-of-the-art algorithms on various offline multi-agent benchmarks.
Key words and phrases:
Offline Multi-Agent Reinforcement Learning; Value Factorization
doi:
UOYL1486
1. Introduction
@@
Reinforcement learning (RL) trains agents to maximize cumulative rewards through interaction with an environment in an online manner; however, such interaction can be time-consuming due to the sample inefficiency of RL and costly in safety-critical environments due to inherent risks levine2020offline; NEURIPS2023_a6f6a5c5. Furthermore, in many practical scenarios, only pre-collected datasets can be leveraged instead of interacting with the environment. These challenges have motivated a shift toward offline RL, which trains agents using pre-collected datasets without requiring interaction kostrikovoffline; wu2019behavior; agarwal2020optimistic; peng2023weighted.
@@
A major challenge of offline RL is that bootstrapping from unseen actions in policy evaluation induces extrapolation error, leading to accumulated errors levine2020offline; fujimoto2021minimalist, due to the inability to generate data. This issue consequently causes overestimation of value functions and adversely affects policy improvement. To address this, a variety of techniques, such as conservative value estimation kumar2020conservative and modeling behavior policy fujimoto2019off, have been introduced. However, these approaches increase algorithmic complexity, hinder reproducibility, and amplify hyperparameter sensitivity. This motivates minimalist approaches, which introduce minimal changes to existing RL algorithms fujimoto2021minimalist; tarasov2024revisiting. One representative example is TD3+BC, which integrates behavior cloning (BC) regularization into an existing online RL algorithm called TD3 fujimoto2021minimalist. This simple approach does not require any additional components and introduces only one more hyperparameter for regularization, but it surprisingly achieves SOTA performance. This line of research, minimalist approaches, has been extensively investigated in single-agent settings fujimoto2021minimalist; tarasov2024revisiting, but it has barely been explored in multi-agent settings, which entail more algorithmic complexity due to the presence of multiple agents.
@@
To address the aforementioned limitations, we propose a simple yet effective technique for offline multi-agent RL: Behavior Cloning regularization with Critic Clipping (B3C). The proposed method alleviates overestimation by clipping the critic value during policy evaluation, using the maximum return from the given dataset, in addition to BC regularization. Critic clipping enables the use of a higher weight for the RL objective relative to BC regularization, resulting in improved performance. Furthermore, we integrate B3C with an existing online multi-agent RL algorithm called FACMAC peng2021facmac, which utilizes factored critics and value factorization techniques, resulting in FACMAC+B3C. Notably, we provide a design choice for value factorization in the offline setting, empirically demonstrating that non-monotonic factorization performs better than monotonic and linear factorization, which contrasts with findings in the online setting peng2021facmac.
@@
Despite its simplicity, the effectiveness of FACMAC+B3C is empirically demonstrated across a variety of offline multi-agent RL benchmarks, including multi-agent Mujoco and particle environments, with various types of dataset. We provide a thorough analysis demonstrating how the proposed method outperforms the baselines in terms of performance. Our main contributions are:
$\bullet$ We address the over-regularization problem in offline multi-agent RL and propose a minimalist framework that achieves SOTA performance with minimal modification to existing algorithms.
$\bullet$ We conduct an empirical study of value factorization in offline settings, an aspect that has been understudied.
$\bullet$ We conduct extensive experiments across seven environments and forty-two datasets, covering both fully and partially observable settings.
2. Background and Related Works
2.1. Multi-Agent RL
@@
We consider fully cooperative multi-agent tasks, which can be modeled as an N-agent decentralized partially observable MDP (Dec-POMDP) oliehoek2012decentralized; dibangoye2018learning; amato2013decentralized. At each timestep $t$, each agent selects its own action, $a_{t}^{i}$, based on its local information such as the partial observation $o_{t}^{i}$. This forms a joint action $\bm{a}_{t}$ yielding the next state $s_{t+1}$, the joint observations $\bm{o}_{t+1}=(o^{1}_{t+1},\cdots,o^{N}_{t+1})$, and a shared reward $r_{t}$. The goal is to find the optimal joint policy that maximizes the team return, $\mathbb{E}[\sum_{l=0}^{\infty}\gamma^{l}r_{l}]$. For this, recent multi-agent RL algorithms adopt a framework of centralized training with decentralized execution (CTDE) wang2020dop; zhang2021fop; jeon2022maser; kim2023variational, where individual policies execute actions based on local information, e.g., Agent $i$’s action-observation history $\tau^{i}$, but are trained with global information such as the state.
One of the key challenges in Dec-POMDPs is to correctly assign credit to each agent for the team reward. For this, value factorization methods have been actively studied sunehag2017value; rashid2020monotonic; son2019qtran; zhang2021fop; wang2021towards; marchesini2025stateful. Value factorization methods approximate the joint $Q$-function as a function of individual agent-wise value functions, thereby enabling effective credit assignment among agents. Linear factorization, such as Value Decomposition Networks (VDN) sunehag2017value, assumes additive decomposition of individual $Q^{i}$ values, while monotonic factorization in QMIX rashid2020monotonic, enforces a monotonic mixing constraint to ensure consistency between local and global optima. Such methods have shown promising results in online multi-agent RL, but their applicability to offline settings remains unexplored.
FACMAC peng2021facmac is a representative multi-agent RL algorithm that trains a centralized but factored critic, which is decomposed into a nonlinear combination of individual critics, and uses it to train decentralized deterministic policies.The key idea is factorizing the centralized critic into individual critics via a non-linear function, where the centralized critic is decomposed as
$Q_{jt}(s,\bm{\tau},\textbf{a})=f_{mixer}(s,Q^{1}(\tau^{1},a^{1}),\cdots,Q^{N}(\tau^{N},a^{N})),$ (1)
@@
where $f_{mixer}$ is a mixer network that encodes state information and learns the weights of individual critics, and $Q^{i}(\tau^{i},a^{i})$ is the individual critic of Agent $i$. FACMAC considers two value factorization methods: (a) monotonic factorization (mono), which constrains the parameters of $f_{mixer}$ to enforce $\partial Q_{jt}/\partial Q^{i}\geq 0$, as in QMIX rashid2020monotonic; and (b) non-monotonic factorization (non-mono), which removes the constraint from (a) to increase representational capacity. In addition to these two methods (c) linear factorization sunehag2017value (vdn) was also considered. It was shown that non-mono and mono approaches are complementary: non-mono performs better on tasks requiring greater representational capacity, while mono outperforms on others. However, vdn underperforms both and performs drastically worse, particularly in multi-agent Mujoco environment.
@@
$\mathcal{L}_{\text{FACMAC}}$ $(\bm{\theta_{Q}})=\mathbb{E}_{\mathcal{D}}\left[\Big(y^{jt}-Q_{jt}(s,\bm{\tau},\bm{a};\bm{\theta_{Q}})\Big)^{2}\right],\penalty 10000\ \mbox{where}$
$y^{jt}=r+\gamma Q_{jt}(s^{\prime},\bm{\tau^{\prime}},\bm{\pi}(\bm{\tau^{\prime}});\bm{\theta_{Q}}^{-}),$ (2)
$\mathcal{D}$ is the replay buffer, $\bm{\pi}$ is the deterministic joint policy, and $\bm{\theta_{Q}^{-}}$ is the parameter of the target critic networks. Based on this centralized critic, the decentralized policies, parameterized by $\bm{\theta}_{\bm{\pi}}=\{\theta_{\pi}^{i}\}_{i=1}^{N}$, are trained using the deterministic policy gradient silver2014deterministic where the loss function is written as $\mathcal{L}_{\text{FACMAC}}(\bm{\theta}_{\bm{\pi}})=$
@@
where $\theta_{\pi}^{i}$ is Agent $i$’s actor parameter. A main benefit of FACMAC over prior works such as MADDPG lowe2017multi is leveraging non-linear factorization. However, in offline settings, value factorization techniques remain underexplored—recent research still relies on linear factorization yang2021believe; wang2024offline or omits it by using a non-factored critic. In this paper, to investigate value factorization for offline settings, we adopt FACMAC to build upon the recent achievements.
2.2. Offline RL
Offline RL aims to learn a policy that maximizes the expected return from a fixed dataset $\mathcal{D}$, consisting of trajectories generated by arbitrary behavior policies, without additional environment interactions kumar2020conservative; kostrikov2022offline; kim2024adaptive. This inability to generate additional data introduces a major challenge in offline RL—extrapolation error in policy evaluation, i.e., the target value in the Bellman equation becomes inaccurate if actions from the learning policy are not included in the dataset. This extrapolation error often leads to overestimation and degrades performance. To address this challenge, several approaches have been proposed, such as behavior-constrained policy optimization, which regularizes the policy to stay close to the behavior policy wu2019behavior; fujimoto2021minimalist; wu2022supported; zhang2023constrained, and conservative value estimation kumar2020conservative; kostrikov2021offline, which lower-bounds the value function by penalizing the values of unseen actions. These approaches have demonstrated effectiveness; however, they often require significant modifications or additional components to existing RL algorithms, such as generative models. These additions introduce more hyperparameters, which can affect performance and impact stability and reproducibility. This necessitates an offline RL algorithm with minimal changes from the current online RL algorithms. As an example, fujimoto2021minimalist proposed TD3 with Behavior Cloning (TD3+BC), which adds a BC loss function as a regularizer on top of the TD3 loss fujimoto2018addressing to update the policy. This encourages the action generated by the learning policy to be the same as the action in the dataset. The policy of TD3+BC is updated to minimize the following loss function.
@@ 2.2. Offline RL
where $\bm{w}=\frac{\alpha}{\frac{1}{N}\sum|Q(s,\pi(s))|}$ (4)
@@
where $\alpha$ is a hyperparameter that controls the balance between RL and BC. TD3+BC has been widely used due to its simplicity and performance comparable to SOTA algorithms kim2024decision.
2.3. Offline Multi-Agent RL
@@
A natural approach to offline multi-agent RL is to extend the previously mentioned single-agent offline RL to the multi-agent setting. However, naively extending these methods has been shown to be insufficient pan2022plan; shao2024counterfactual. For example, naive conservative value estimation for the joint policy can lead to excessively large penalties, significantly degrading performance as the number of agents increases shao2024counterfactual. Furthermore, the problem of extrapolation error worsens as the size of the action space grows with the number of agents yang2021believe, making it essential to balance conservatism and regularization with extrapolation error. To mitigate this problem, CFCQL shao2024counterfactual introduces counterfactual regularization to each agent individually to alleviate excessive conservatism. OMAR pan2022plan introduces zeroth-order optimization to address policies getting stuck in poor local optima under conservative value function training. Note that both OMAR and CFCQL are based on conservative value estimation, which penalizes value functions for unseen action spaces, thereby forcing them to underestimate. OMIGA wang2024offline introduces a framework that converts global regularization into implicit local-level regularization using linear value decomposition, while employing in-sampling techniques to avoid querying unseen actions. MADIFF zhu2023madiff presents a diffusion-based generative approach for multi-agent systems, enabling decentralized policy development through effective teammate modeling.
@@
Although the aforementioned offline multi-agent RL algorithms perform reasonably, they still require significant modifications from existing online multi-agent RL algorithms, as described in single-agent settings in Sec. 2.2, with even more adjustments needed due to the presence of multiple agents. Additionally, they have not fully leveraged the currently existing value factorization techniques, which are central to the recent success of multi-agent RL. Therefore, inspired by fujimoto2021minimalist, we aim to develop a minimalist approach that minimally modifies an existing multi-agent RL algorithm, focusing on simplicity and practicality while incorporating existing multi-agent RL techniques.
3. Methodology
To achieve the aforementioned goal, we focus on a BC regularization-based approach. In multi-agent settings, the presence of multiple agents amplifies the issue of overestimation, as mentioned in Sec.2.3. Consequently, the BC regularization-based approach is susceptible to either over-regularization or unstable critic learning. This occurs because the joint action space grows exponentially with the number of agents, increasing the likelihood of encountering unseen actions during policy evaluation. As a result, critic updates become unstable and prone to divergence, which in turn encourages excessive reliance on the BC term to stabilize training, leading to over-regularization and performance degradation when the dataset quality is limited.
@@ 3. Methodology
To address these challenges, we propose a simple yet effective regularization method called Behavior Cloning with Critic Clipping (B3C), which prevents over-regularization while ensuring stability, allowing a focus on the RL objective and thereby significantly improving performance. Additionally, we explore the application of value factorization techniques with B3C, which remains understudied.
@@
Figure 1. Test return as a function of dataset quality (average dataset return) in CN and World, based on shao2024counterfactual. Proximity to $y=x$ reflects strong data dependence. MA-TD3+BC remains close to dataset-level performance, indicating over-regularization by behavior cloning.
3.1. Revisiting Behavior Cloning: Over-regularization
We begin by revisiting BC regularization in offline multi-agent RL. Building on the success of BC regularization in the single-agent setting, recent studies on offline multi-agent RL pan2022plan; shao2024counterfactual have introduced an extension of TD3+BC, called MA-TD3+BC, as a baseline by adapting Eq. 2.2 to the multi-agent setting. Although MA-TD3+BC achieves reasonable performance, it is less effective than the proposed methods, particularly when applied to low-quality datasets. Fig. 1 illustrates the test returns of OMAR pan2022plan, CFCQL shao2024counterfactual, and MA-TD3+BC with respect to the average return of the dataset, reported in shao2024counterfactual. MA-TD3+BC achieves performance close to dataset-level quality, but with a small margin. In contrast, OMAR and CFCQL show significant improvements over dataset quality, particularly for low-quality datasets. This observation, that MA-TD3+BC heavily depends on data quality, indicates that it is over-regularized by behavior cloning.